OpenStudy (lyricscompleteme):

Which line is not perpendicular to the line 4x − 3y = 2?

OpenStudy (lyricscompleteme):

3x + 4y = 2 3x − 4y = 2 -3x − 4y = 5 3x + 4y = -8 Are the choices..

OpenStudy (anonymous):

if the slope of line 1 is m1 and slope of line 2 is m2, then they are perpendicular when m1 * m2=-1

OpenStudy (lyricscompleteme):

I got it it was B. Thank you! c:

OpenStudy (anonymous):

it should be D

OpenStudy (anonymous):

check your answer again please, and use the formula i have provided to you